OpenStudy (anonymous):

Find the exact value for sin((3pi/4) + (5pi/6))

OpenStudy (freckles):

use sum identity for sin

OpenStudy (freckles):

that is sin(x+y)=sin(x)cos(y)+sin(y)cos(x)

OpenStudy (freckles):

then you would use the unit circle to evaluate sin(3pi/4) sin(5pi/6) cos(3pi/4) cos(5pi/6)

OpenStudy (freckles):

plug in and follow the order of operations

OpenStudy (jhannybean):

When using the unit circle to find your values, find your reference angles.

OpenStudy (anonymous):

So would that be like this - > sin(3pi/4 + 5pi/6) = sin(3pi/4)cos( 5pi/6) + sin(5pi/6)cos(3pi/4) //Sorry for the slow reply

OpenStudy (freckles):

oh yes

OpenStudy (freckles):

sorry I have multiple windows open :p

OpenStudy (anonymous):

It's fine, and then like you said I'd just follow order of operations?

OpenStudy (freckles):

yep once you used the unit circle to evaluate all of those thingys

OpenStudy (jhannybean):



OpenStudy (anonymous):

Um, well I feel dumb, but how do you do order of operations with sin/cos? Like do I just solve for the sin / cos values and then add them and compare? Or is it more involved than that? Also, Jhanny how do the reference angles figure into find the exact values? //I've been kinda lacking in my Trig studies.

OpenStudy (freckles):

@mutedwolf you have to used the unit circle to evaluate sin(3pi/4) sin(5pi/6) cos(3pi/4) cos(5pi/6) then use order of operations

OpenStudy (freckles):

for example sin(3pi/4)=sqrt(2)/2

OpenStudy (anonymous):

Ohhh, okay, that makes more sense.

OpenStudy (jhannybean):

Yeah....that's why I drew you a unit circle, lol.

OpenStudy (anonymous):

Ah, lol, okay. So would the evaluated equation look like this? -> (sqrt (2)/2 + sqrt 1/2) = (sqrt (2)/2 * -sqrt 3/2) + sqrt 1/2 * -sqrt 2/2)

OpenStudy (freckles):

to the left what you have is off but to the right it looks good

OpenStudy (jhannybean):

You were asking about the reference angles, totally skipped over that post. Sorry. The reference angles tell you the x an y values in the other 3 quadrants. Because the reference angles are all positive, it's easier to use them than calculating what the x and y components in the other quadrants will be - just look at quadrant one for all your answers :)

OpenStudy (jhannybean):

The only thing that changes is their sign

OpenStudy (freckles):

\[\sin(\frac{3\pi}{4}+\frac{5\pi}{6})=\frac{\sqrt{2}}{2} \frac{-\sqrt{3}}{2}+\frac{1}{2} \frac{-\sqrt{2}}{2}\] just wanted to change what you had above because it was icky to me :p

OpenStudy (anonymous):

Ohh, I was getting mixed up about left geez, thought you meant one of my unit circle values was off.

OpenStudy (freckles):

no I just want you to know sin(x+y) doesn't equal sin(x)+sin(y) and that is what it look like what you were doing on the left

OpenStudy (freckles):

\[\text{ so you actually found what } \sin(\frac{3\pi}{4}+\frac{5\pi}{6}) \text{ or } \sin(\frac{19}{12} \pi) \text{ equals exactly }\]

OpenStudy (freckles):

I would multiply a little and combine the fractions though

OpenStudy (anonymous):

Thanks for the help!